40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 228 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  5401   Wed Sep 14 01:19:20 2011 AnamariaConfigurationLSC3f PD Install in Progress

I have reconfigured the refl beam path on the AP table to include REFL33 and REFL165. Would be done if we hadn't prepared P BSs instead of S, which required some serious digging to find two others. And if someone hadn't stolen our two 3m SMA cables that Keiko and I made on our previous visit and I had left with the 3f PDs. I don't expect them to reappear but if they do, it would be grand.

Note: Refl beam from ifo looks a bit high, ~1cm on the lens 20'' from output port. Not sure what that means about ifo alignment change, I've left it as is. When we know we have a good alignment, we should be able to easily realign the beam path if necessary. If it remains the same, we might want to change the lens height.

Done:

1) REFL11 and REFL55 are now hooked up and aligned in a low power beam. (I set the power as low as I could by eye to not risk burning the PDs during alignment)

2) The required BSs and REFL33 and REFL165 are in place, powered.

3) I have set them in a configuration such that the beam is the same distance from the main beam, to adjust beam size easily for all 4.

4) Camera has been moved from main beam to behind a steering mirror, ND filters removed, centered on camera.

To Do:

1) Find one more longish SMA cable.

2) Align beam on REFL33 and REFL165.

3) Check beam size carefully. (I get a plateau on the scope, and I can "hide" the beam on the PD, but it could be better. The path has become longer by ~5-8inches.)

4) Adjust power.

5) Redo layout diagram, post in wiki.

  5400   Wed Sep 14 01:17:51 2011 kiwamuUpdateLSCmeasurement of spot position on Y arm

The spot positions on ITMY and ETMY were measured using the LOCKIN modules in C1ASS when the Y arm stayed locked.

The beam was successfully aligned such that it hits the center of the ETMY mirror.

However on the other hand the angle of the beam is pitching and it's going upward as the beam propagates to ETMY.

 

/***** RESULTS ******/

Here is a summary of the measurement :

  amount of off-centering [mm]
ETMY_PIT  0.0032
ETMY_YAW  -0.13
ITMY_PIT  -12
ETMY_YAW  0.91

  

Also a cartoon is shown below.

The scale is not quite true, but at least it gives you a 3D information of how the beam is pointing down to the Y arm.

 misposition.png


  /***** MEASUREMENT *****/

 In order to measure the spot positions the standard technique, namely A2L, was used.

Since the C1ASS model was made for doing the A2L measurements on each arm cavity, the LOCKIN modules in C1ASS were used.

First the Y arm was locked with AS55 (#5398), and then the C1ASS was activated by calling some scripts from C1ASS_QPDs.adl.

In order to calibrate the signals from LOCKINs, an intentional coil imbalance was introduce.

This is the same calibration technique as Valera explained before (#4355) for measurement of the MC spot positions.

Quote from #5398

The Y arm has been locked with AS55.

A next thing is to check the spot positions on the ETMY and ITMY mirrors so that we can evaluate the recent beam pointing.

 

  5399   Tue Sep 13 23:08:51 2011 KeikoUpdateLSC3f demodulation board check

Keiko, Jamie , Kiwamu

The I and Q orthogonalities of REFL33 and 165 demodulation board were measured by "orthogonality.py"  Python package scipy were addied on Pianosa to run this code. Please note that "orthogonality.py" can be run only on Pianosa.  

The results were:

REFL165

ABS = 1.070274, PHASE = -81.802479 [deg]

if you wanna change epics values according to this result, just copy and execute the following commands

ezcawrite C1:LSC-REFL165_Q_GAIN 0.934340 && ezcawrite C1:LSC-REFL165_PHASE_D -81.802479

- - - - - - - - - - - - - - - - - -

REFL33

ABS = 1.016008 PHASE = -89.618724 [deg]

if you wanna change epics values according to this result, just copy and execute the following commands

ezcawrite C1:LSC-REFL33_Q_GAIN 0.984244 && ezcawrite C1:LSC-REFL33_PHASE_D -89.618724


Fig.1 and 2  are the resulting plots for 33 and 165 MHz demod baoards, respectively.You should look at the 3Hz in x axis, as the demodulated signal frequency was set as 3 Hz.REFL33-modified.png

Fig. 1 REFL33 I and Q orthogonality at 3 Hz.

REFL165-modified.png

 

Fig. 2 REFL165 I and Q orthogonality at 3 Hz.

 

 

 

  5398   Tue Sep 13 19:31:09 2011 kiwamuUpdateLSCY arm locking prep

The Y arm has been locked with AS55.

A next thing is to check the spot positions on the ETMY and ITMY mirrors so that we can evaluate the recent beam pointing.

 

- - - parameter settings - - -

C1:LSC-YARM_GAIN = -0.03

AS55 demod phase = 0.2

WF gains = 21 dB

C1:LSC-TRY_OUT = 0.57 (maximized by steering PZT2)

Quote from #5392

This was the reason why I had to do the f2a filter redesign. Tomorrow we can try locking the Y arm.

 

Attachment 1: Yarmlocked.png
Yarmlocked.png
  5397   Tue Sep 13 19:12:53 2011 SureshUpdateComputer Scripts / Programschanges to the c1ioo, c1mcs and c1rfm

[Suresh, Jamie]

The new versions of these three models have been committed to the svn.  I copy below the svn log I wrote over there.

I edited three models,namely c1ioo, c1mcs and c1rfm to bring four channels from C1MCS model into the
 C1IOO model through the RFM.

We plan to use the MC2-Trans-QPD as a third sensor in addition to the WFS1 and WFS2 in the ASC system so that we can sense all the six degrees of freedom of the MC.  However the MC2-Trans_QPD is a new incarnation of the old MC2 Oplev QPD and its four segments are are picked up in the c1mcs model running on the C1SUS machine.  Since we need these    in the C1IOO machine we    had to port these signals via the RFM.  The    changes    I made to these    models are described below:
1) c1rfm.mdl -->  I added several cds parts to carry these four    channels from c1mcs to c1ioo via c1rfm

2) c1mcs.adl -->  In the MC2 block of c1mcs I cut out the structure which converts the four segments of the  QPD in the pitch, yaw and sum signals and copied it to the c1ioo model.     In the    c1mcs I  grounded the input of the oplev signals just as they are in    the MC1    and MC2    suspension blocks. I then routed the ADC channels associated with the oplev QPD signals to the rfm model.

3) c1ioo.mdl --> I have made extensive changes to the c1ioo model to include the third sensor in the WFS feedback loop.  The six lockins in the model are used to excite the MC mirrors and the signals from WFS and QPD can be demodulated to determine the input or output matrices in the WFS feedback.  A sensor_demod_array permits us to select the signals we wish to demodulate. The MCL signal, used in the MC_ASS and A2L measurements, is retained as one of the signals.  The MC_TRANS_SUM was used several in  scripts and to preserve the continuity with this channel name we inserted three EPICS parts to generate a slow channel by the same name from this model.

Since the MC_TRANS_SUM is used in the MC-autolocker script we checked to make sure that the script is running after we recompiled all the three models and restarted them.  It is all fine.

  5396   Tue Sep 13 19:04:58 2011 SureshUpdateComputer Scripts / ProgramsRestarted Frame builder several times while compiling c1ioo, c1mcs and c1rfm

I restarted the frame builder at the following times

 

Tue Sep 13 14:53:49 PDT 2011

 

Tue Sep 13 16:46:32 PDT 2011

 

Tue Sep 13 17:24:16 PDT 2011

  5395   Tue Sep 13 16:57:44 2011 steveUpdateLSCITMX oplev **

Quote:

I lined up the Y Arm for locking and then centered the oplevs for ETMY and ITMY.

* The ITMY OL has still got the old style laser. Steve, pleaes swap this one for a HeNe. Also the optical layout seems strange: there are two copies of the laser beam going into the chamber (??). Also, the QPD transimpedance needs to be increase by a factor of ~10. We're only getting ~500 counts per quadrant. Its worth it for someone to re-examine the whole ITMY OL beam layout.

* The ETMY OL beam was coming out but clipping on the mount for the ETMY OL HeNe. This indicates a failure on our part to do the ETMY closeout alignment properly. In fact, I get the feeling from looking around that we overlooked aligning the OL and IPPOS/ANG beams this time. If we're unlucky this could cause us to vent again. I undid part of the laser mount and changed the height on the receiving mirror to get the beam back onto the QPD.

I noticed that there is significant green light now getting into some of the IR PDs; beacuse of this there are weird offsets in the TRY QPD and perhaps elsewhere. We had better purchase some filters to tape over the front of the sensitive IR sensors to prevent the couplling from the green laser.

* There is a beam on IPPOS, but its too big for the detector (this has always been the case). We need to put a 2" lens with a weak focusing power on this path so as to halve the beam size on the detector. Right now its clipping and misleading. There is also a 0.9V offset on the SUM signal. I'm not sure if this readout is working at all.

* I couldn't find any beam on IPANG at all. Not sure what's changed since Kiwamu saw it.

 Old Coherent diode laser was replaced by Uniphase HE/NE 1125P at ITMX.

Newport 10B20NC.1 broadband beam sampler set up as beamspitter to dump 90% of the light into beam trap. Beam in to test mass  0.7 mW ,  returning to qpd 0.07 mW  = ~ 3,700 counts

f 1.5 m lens placed into ingoing beam path to reduce the spot size on the qpd.         ITMY oplev will be done in this manner tomorrow.

 

* IP-ANG is lost with the Piezo Jena PS

* IP-POS cable have to be found at the LSC rack

 

  5394   Tue Sep 13 15:00:25 2011 KeikoUpdateLSC3f demodulation board check

Whitening filters for the REFL33 & 165 demodulated channels were measured and confirmed that they are working. They can be turned on and off by un-white filter switches on the MEDM screen because they are properly linked. The measured filter responses are showen below. (Sorry, apparentyl the thumbnails are not shown here. Please click the attachments.) 

WF33.pdf

WF165.pdf

Attachments: (top) Whitening filter for REFL33 demodulation board. (bottom) Whitening filter response for REFL 165 demodulation board.

  5393   Tue Sep 13 09:16:00 2011 SureshUpdateIOOMC spots recentered and input beam aligned

The shift of MC2 which Rana noted caused the beam spots on the MC mirrors to decenter. I used the mcassUp and mcassOn scripts and checked the output of the C1IOO lockins to get the spot positions.  I first tried to realign just the MC2 to recenter the spots.  But this was not sufficient.  I then worked on the pitch of all three optics since it is easier to align.   By the time this was done the offset in yaw also reduced, probably due to cross coupling between pitch and yaw in the coils.  At the end of the process I obtained all decentering around 1.5mm or less, then I went over to adjust the MC2TransQPD beam path so that we center the spot on the QPD.  This action shifted the stack,  I had to iterate this two more times before the successive corrections grew sufficiently small.  I think it may shift again if we touch the chamber (the image of MC2Face is still inverted).

The new spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    1.3212   -0.8415    0.6795   -1.4292   -0.3574   -1.5208

 

 mcdecenter20110912_1.png    mcdecenter20110912_2.png      

 

- Further improvement of beam centering can be done but first I would like to be sure that the MC is stable.  The MC2Trans light is centered on the QPD as a reference.

 

  5392   Tue Sep 13 03:18:14 2011 kiwamuUpdateLSCY arm locking prep

(Preparation of Y arm locking)

(A) The f2a filters were newly designed and applied to ETMY (see the attachment)

(B) Once the Y arm is aligned such that the TEM00 mode flashes, the transmitted light is visible on the ETMYT CCD camera.

(C) With the newly installed resonant EOM circuit the PDH signal from AS55 looks healthy.

 

(some notes)

(A) To design the f2a filters there is a handy python script called "F2A_LOCKIN.py" in /scripts/SUS.

The script measures the coil imbalance at high frequency and low frequency using a LOCKIN module and then gives us the information about the imbalance.

The script hasn't yet been completed, so it doesn't return the intuitive answers but returns something non-intuitive. I will modify it.

 

(B)  To see the transmitted light from the Y arm I was going to align the CCD camera on the Y end table.

However I found that once the green light is blocked, the transmitted light can be visible on the camera without any re-alignment.

Therefore I haven't rearranged anything on the Y end table, but I just blocked the green light.

Perhaps we still need to align the photo diodes for the transmitted light.

 

(C) While Suresh was working on MC, I looked at the signal from AS55 with all the optics misaligned except for ITMY, ETMY and BS.

The signal from the Y arm looked very PDH signal, and the demodulation phase seemed to be about 45 deg to maximize the I signal.

I tried locking it by feeding the signal back to ETMY but failed due to a too much POS to angle coupling in the ETMY actuators.

I was momentarily able to capture a higher order mode with a negative gain in LSC-YARM_GAIN, but it was quite difficult to keep it locked.

This was because once I increased the gain to make it stable, the angle instability became more significant and lost the lock immediately.

This was the reason why I had to do the f2a filter redesign. Tomorrow we can try locking the Y arm.

Attachment 1: image.png
image.png
  5391   Mon Sep 12 23:54:10 2011 kiwamuUpdateIOOEOM resonant box installed

[Mirko / Kiwamu]

 The resonant box has been installed together with a 3 dB attenuator.

The demodulation phase of the MC lock was readjusted and the MC is now happily locked.

 

(Background)

We needed more modulation depth on each modulation frequency and so for the reason we installed the resonant box to amplify the signal levels.

Since the resonant box isn't impedance matched well, the box creates some amount of the RF reflections (#5339).

In order to reduce somewhat of the RF reflection we decided to put a 3 dB attenuator in between the generation box and the resonant box.

 

(what we did)

 + attached the resonant box directly to the EOM input with a short SMA connector.

 + put stacked black plates underneath the resonant box to support the wight of the box and to relief the strain on the cable between the EOM and the box.

 + put a 3 dB attenuator just after the RF power combiner to reduce RF reflections.

 + readjusted the demodulation phase of the MC lock.

 

(Adjustment of MC demodulation phase)

 The demodulation phase was readjusted by adding more cable length in the local oscillator line.

After some iterations an additional cable length of about 30 cm was inserted to maximize the Q-phase signal.

So for the MC lock we are using the Q signal, which is the same as it had been before.

 

 Before the installation of the resonant box, the amplitude of the MC PDH signal was measured in the demodulation board's monitor pins.

The amplitude was about 500 mV in peak-peak (see the attached pictures of the I-Q projection in an oscilloscope). Then after the installation the amplitude decreased to 400 mV in peak-peak.

Therefore the amplitude of the PDH signal decreased by 20 %, which is not as bad as I expected since the previous measurement indicated 40 % reduction (#2586).

 

Attachment 1: EOM.png
EOM.png
Attachment 2: before.png
before.png
Attachment 3: after.png
after.png
  5390   Mon Sep 12 23:45:14 2011 SureshConfigurationComputer Scripts / ProgramsStripTool does not scale properly on Pianosa

When I run StripTool on Pianosa, I get the following message

==== StripTool Xt Warning Handler ====
warning:         Axis: minVal is greater than or equal to maxVal

 

And the y-axis scale reverts to 0 -100 regardless of what ever I set in the controls panel

I ssh'ed into rosalba and ran striptool from there and did not face this problem.  So I think pianosa has a problem with Striptool.

  5389   Mon Sep 12 18:45:04 2011 AnamariaConfigurationLSCAP table current layout

Before we install the REFL 3f PDs I made a drawing of the current table layout, since there has been no update lately. Once I've incorporated the two extra PDs (now seen sitting bottom left), I will update the drawing and post in the wiki as well.

Attachment 1: 40mAPtable.pdf
40mAPtable.pdf
  5388   Mon Sep 12 18:40:35 2011 KeikoUpdateLSC3f demodulation board check

LSC-REFL33-I-IN1(IN2, OUT) and LSC-REFL165-Q-IN1(IN2,OUT) channels are back!

We disconnected and connected again the AA filters then the channels are fixed. Apparently the AA filters just before the digital world were somhow charged and not working... Thank you Kiwamu!

Quote:

Wait. I am checking the whitening filters of the 33 and 165 demodulation boards.

Also, LSC-REFL33-I-IN1(IN2, OUT) and LSC-REFL165-Q-IN1(IN2,OUT) channels may not be working??

 

  5387   Mon Sep 12 16:27:01 2011 KeikoUpdateLSC3f demodulation board check

Wait. I am checking the whitening filters of the 33 and 165 demodulation boards.

Also, LSC-REFL33-I-IN1(IN2, OUT) and LSC-REFL165-Q-IN1(IN2,OUT) channels may not be working??

 

Quote:

I also quickly checked the orthogonality of the demodulation board for REFL33 and REFL165 using function generators and oscilloscope. I checked the frequencies at 1,10,100,1K,10KHz of the demodulated signals. They are fine and ready for 3f signal extraction.

 

  5386   Mon Sep 12 13:24:07 2011 KeikoUpdateLSC3f demodulation board check

I also quickly checked the orthogonality of the demodulation board for REFL33 and REFL165 using function generators and oscilloscope. I checked the frequencies at 1,10,100,1K,10KHz of the demodulated signals. They are fine and ready for 3f signal extraction.

  5385   Sun Sep 11 22:36:32 2011 KeikoUpdateLSC3f demodulation board check

Filters at the RF inputs of REFL33 and REFL165 demodulation boards were measured again. The filters will be totally fine for 33MHz and 165MHz.

Last time I forgot to calibrate the cable lengths, therefore the phase delay of the measurement included the cable lengths. This time the measurements were done for REFL33 and REFL165 demod board with calibration. As the cable lengths were calibrated, the shown plots (Fig.1 and Fig.2) do not include the phase delay dues to measurement cables. Please note that the x-axis is in linear. The phase delays of both boards seems to be not too steep (it will not affect anyway, as Kiwamu pointed out in his comment on the previous post). You can see that the two filters do not filter 33MHz and 165MHz component out.

REFL33.png

Fig.1 A response of a filter which is placed just after the RF input of the demodulation board for REFL33. X-axis is shown in linear (~50MHz).

REFL165.png

Fig.2 A response of a filter which is placed just after the RF input of the demodulation board for REFL165.

 

Quote:

There is a LP filter just after the RF input of an demodulation board (its schematic can be found as D990511-00-C on DCC). I have checked if the 3f freq, 33MHz, can pass  this filter. The filter TF from the RF input to RF monitor (the filter is between the input and monitor) on REFL33 demo-board was measured as shown in Fig. 1. At 33MHz, the magnitude is still flat and OK, but the phase is quite steep. I am going to consider if it is ok for the PDH method or not.

REFL33-input-filter.png

 Fig. 1 Transfer function from the RF input to RF monitor on the REFL33 demodulation board. At 33MHz, a very steep phase is applied on the input signal.

Quote:

 To check the demodulation boards for REFL33 and REFL165, a long cable from ETMY (SUS-ETMY-SDCOIL-EXT monitor) is pulled to the rack on Y side.

(1) A filter just after the RF input and (2) transfer function from the RF input to the demodulated signal will be checked for the two 3f demod boards to confirm that they are appropriate for 33 and 165 MHz.

 

 

  5384   Sun Sep 11 15:19:39 2011 kiwamuUpdateLSCphase delay in RF signals

A comment :

Since the LSC RFPD have a long cable of more than 6 m, which rotates a 33 MHz signal by more than 360 deg, so the delay has always existed in everywhere.

The circuit you measured is a part of the delay existing in the LSC system, but of course it's not a problem as you said.

In principle a delay changes only the demodulation phase. That's how we treat them.

RA: Actually, the issue is not the delay, but instead the dispersion. Is there a problem if we have too much dispersion from the RF filter?

 

  5383   Sat Sep 10 20:30:01 2011 ranaUpdateIOOMC trans re-aligned / MC2 shifted mysteriously / MC2 re-aligned

MCdecenter.png

I re-aligned the beam onto the MC TRANS QPD since Kiwamu had centered the spots on the mirrors. However, I then inspected the MC2F camera. After coming back into the control room I noticed that the MC transmission had gone down by 50% and that the MC2 OSEMs showed a large step. My guess is that somehow the opening and closing of the can shifted the suspension. So I adjusted the MC2 alignment biases to recover the transmitted power (its now ~50000 instead of the ~33000 from Friday).

  5382   Sat Sep 10 19:45:29 2011 JenneUpdateLSCY Arm work

Quote:

 * ITMY OL: Also the optical layout seems strange: there are two copies of the laser beam going into the chamber (??).

* The ETMY OL beam was coming out but clipping on the mount for the ETMY OL HeNe. This indicates a failure on our part to do the ETMY closeout alignment properly.

 The 2nd beam from this laser is for the SRM's OpLev, so that shouldn't be changed.

For better or worse, we didn't do anything to the ETM OpLevs, because they don't have any in-vac steering optics.  We did however go through and check on all the corner OpLevs.

  5381   Sat Sep 10 19:03:57 2011 ranaUpdateLSCY Arm work

I lined up the Y Arm for locking and then centered the oplevs for ETMY and ITMY.

* The ITMY OL has still got the old style laser. Steve, pleaes swap this one for a HeNe. Also the optical layout seems strange: there are two copies of the laser beam going into the chamber (??). Also, the QPD transimpedance needs to be increase by a factor of ~10. We're only getting ~500 counts per quadrant. Its worth it for someone to re-examine the whole ITMY OL beam layout.

* The ETMY OL beam was coming out but clipping on the mount for the ETMY OL HeNe. This indicates a failure on our part to do the ETMY closeout alignment properly. In fact, I get the feeling from looking around that we overlooked aligning the OL and IPPOS/ANG beams this time. If we're unlucky this could cause us to vent again. I undid part of the laser mount and changed the height on the receiving mirror to get the beam back onto the QPD.

I noticed that there is significant green light now getting into some of the IR PDs; beacuse of this there are weird offsets in the TRY QPD and perhaps elsewhere. We had better purchase some filters to tape over the front of the sensitive IR sensors to prevent the couplling from the green laser.

* There is a beam on IPPOS, but its too big for the detector (this has always been the case). We need to put a 2" lens with a weak focusing power on this path so as to halve the beam size on the detector. Right now its clipping and misleading. There is also a 0.9V offset on the SUM signal. I'm not sure if this readout is working at all.

* I couldn't find any beam on IPANG at all. Not sure what's changed since Kiwamu saw it.

  5380   Sat Sep 10 18:57:52 2011 KeikoUpdateLSC3f demodulation board check

The phase delay due to the RF input filter on the demodulation board will not bother the resulting PDH signals.

I quickly calculated the below question (see the blue sentence in the quote below). I applied an arbitrary phase delay (theta) due to the filter I measured, on the detected RF signal by the photo detector. Then the filtered RF signal is multiplied by cos(omega_m) then filter the higher (2 omega_m) freqency as the usual mixing operation for the PDH signal. As a result, the I signal is delayed by cos(theta) and the Q signal is delayed by sin(theta). Therefore the resulting signals and its orthogonalitity is kept ok. From the sideband point of view, theta is applied on both upper and lower and seems to make the unbalance, however, as it is like a fixed phase offset on both SBs at the modulation frequency, the resulting signals is just multiplied by cos or sin theta for I and Q, respectively. It won't make any strange effect (it is difficult to explain by sentence not using equations!).

Quote:

There is a LP filter just after the RF input of an demodulation board (its schematic can be found as D990511-00-C on DCC). I have checked if the 3f freq, 33MHz, can pass  this filter. The filter TF from the RF input to RF monitor (the filter is between the input and monitor) on REFL33 demo-board was measured as shown in Fig. 1. At 33MHz, the magnitude is still flat and OK, but the phase is quite steep. I am going to consider if it is ok for the PDH method or not.

 Fig. 1 Transfer function from the RF input to RF monitor on the REFL33 demodulation board. At 33MHz, a very steep phase is applied on the input signal.

Quote:

 To check the demodulation boards for REFL33 and REFL165, a long cable from ETMY (SUS-ETMY-SDCOIL-EXT monitor) is pulled to the rack on Y side.

(1) A filter just after the RF input and (2) transfer function from the RF input to the demodulated signal will be checked for the two 3f demod boards to confirm that they are appropriate for 33 and 165 MHz.

 

 

  5379   Sat Sep 10 18:52:45 2011 ranaUpdateComputersconlog getting filled up

One of the reasons that conlog seems so slow lately is that its been writing 100's of MB of .log files every day since early summer. It looks like the people who have been working on the mdl builds have not been properly adjusting the conlog channel lists. When this is not done conlog just gets filled up with non-control channels like OUT, OUTPUT, OUTMON, etc.

Peter Shawhan has supplied us with many scripts in the conlog directory to clean up these bloated files and fix the channel list.

  5378   Sat Sep 10 16:10:42 2011 KeikoUpdateLSC3f demodulation board check

There is a LP filter just after the RF input of an demodulation board (its schematic can be found as D990511-00-C on DCC). I have checked if the 3f freq, 33MHz, can pass  this filter. The filter TF from the RF input to RF monitor (the filter is between the input and monitor) on REFL33 demo-board was measured as shown in Fig. 1. At 33MHz, the magnitude is still flat and OK, but the phase is quite steep. I am going to consider if it is ok for the PDH method or not.

REFL33-input-filter.png

 Fig. 1 Transfer function from the RF input to RF monitor on the REFL33 demodulation board. At 33MHz, a very steep phase is applied on the input signal.

Quote:

 To check the demodulation boards for REFL33 and REFL165, a long cable from ETMY (SUS-ETMY-SDCOIL-EXT monitor) is pulled to the rack on Y side.

(1) A filter just after the RF input and (2) transfer function from the RF input to the demodulated signal will be checked for the two 3f demod boards to confirm that they are appropriate for 33 and 165 MHz.

 

  5377   Sat Sep 10 14:55:28 2011 KeikoUpdateLSC3f demodulation board check

 To check the demodulation boards for REFL33 and REFL165, a long cable from ETMY (SUS-ETMY-SDCOIL-EXT monitor) is pulled to the rack on Y side.

(1) A filter just after the RF input and (2) transfer function from the RF input to the demodulated signal will be checked for the two 3f demod boards to confirm that they are appropriate for 33 and 165 MHz.

  5376   Sat Sep 10 11:07:37 2011 ranaHowToSUSOptical Lever Servo Tuning thoughts

Now that we are in a moderately stable condition, its time to design the optical lever feedback transfer functions. We should think carefully about how to do this optimally.

In the past, the feedback shape was velocity damping from 0-10 Hz, with some additional resonant gain around the pendulum and stack modes. There were some low pass filters above ~30 Hz. These were all hand tuned.

I propose that we should look into designing optimal feedback loops for the oplevs. In principle, we can do this by defining some optimal feedback cost function and then calculate the poles/zeros in matlab.

How to define the cost function (? please add more notes to this entry):

1) The ERROR signal should be reduced. We need to define a weight function for the ERROR signal: C_1(f) = W_1(f) * (ERR(f)^2)

2) The OL QPDs have a finite sensing noise, so there is no sense in suppressing the signal below this level. Need to determine what the sensing noise is.

3) The feedback signal at high frequencies (30 Hz < f < 300 Hz) should be low passed to prevent adding noise to the interferometer via the A2L coupling. It also doesn't help to reduce this below the level of the seismic noise. The cost function on the feedback should be weighted apprpriately given knowledge about the sensing noise of the OL, the seismic noise (including stack), and the interferometer noise (PRC, SRC, MICH, DARM).

4) The servo should be stable: even if there is a negligible effect on the ERROR signal, we would not want to have more than 10 dB of gain peaking around the UGFs.

5) The OL QPDs are dominated by drift of the stack, laser, etc. at some low frequencies. We should make sure the low frequency feedback is high passed appropriately.

6) Minimize transmitted power rms in single arm lock etc.

  5375   Sat Sep 10 02:28:45 2011 kiwamuUpdateSUSfree swinging test in vacuum condition

All the optcs were excited

Sat Sep 10 02:14:11 PDT 2011
999681266

 

  5374   Sat Sep 10 01:36:15 2011 kiwamuUpdateASCinterferometer coarsely aligned

The interferometer was coarsely aligned.

Now spatially overwrapped DRMI and FP arm fringes are visible on the AS camera although the incident beam alignment was done only with PZT2.

All the DC biases were saved so that we can go back to this condition any time.

 

/***** some health checks *******/

 [FINE] IPPOS : it looks okay but the spot on the QPD is a little bit too low by a few mm.

 [NOT GOOD] IPANG : maybe hitting a post or something because the spot is vertically split into two. The spot is too low.

 [FINE] POX/POY/POP : they all are coming out. POP is visible with an IR viewer.

 [FINE] REFL :  no clipping but the beam looks a little bit too low relative to the CCD camera.

 [FINE] AS : no clipping and the spot position on the AS camera looks fine.

 [FINE] Green beams : both X and Y beams are successfully landing onto the PST table without no clipping.

 [FINE] Suspensions : all of them are reasonably quiet without the oplevs, which is good.

  5373   Fri Sep 9 20:52:59 2011 kiwamuUpdateIOOspot positions on MC mirrors (in-vac)

The spot positions on the MC mirrors were readjusted.

All the amount of the off-center became smaller than 2 mm, which meet requirements of the beam clearance on the Faraday.

 

    Feb 26 2011      May 08 2011 Aug 2 2011  Aug 10 2011 (in air) Aug 14 2011 (in air)

Aug 23 2011 (in air)

 Sep 9 2011 [NEW] Sep 9 2011
MC1 pit [mm]   1.6   1.9  1.93 -0.858 -0.2 0.1354  -1.55  0.22
MC2 pit [mm]   6.4   9.0 9.03 -0.844 -0.8 -0.2522  -2.28  -0.34
MC3 pit [mm]   1.4   2.0 2.01 -1.03 -0.1 -0.1383  -1.8 -0.21
MC1 yaw [mm]   -1.5   -1.7 -1.72 -0.847 -1.1 -1.0893  -0.02 -1.15
MC2 yaw [mm]   1.0   0.2 0.178 0.582 0.6 0.7122  3.52 0.07
MC3 yaw [mm]   -1.3   -1.9 -1.87 -1.06 -1.1 -1.5587  -2.145 -1.07

 

 In order to improve the MC1-YAW and MC3-YAW spot positions, the angle of the incident beam has to be shifted by approximately 1/100 rad.

However it turned out to be very difficult to introduce such amount of angle only with the steering mirrors on the PSL table since we have to keep the same translation while changing the angle.

 

Quote from #5371

The obtained spot positions are quite bad and roughly at 2-3 mm level. We have to realign the beam axis and the MC mirrors.

 

  5372   Fri Sep 9 19:15:17 2011 ranaUpdateIOORF Amp for EOM on PSL Table

Quote:

After Steve pointed out the 'deep hoop' issue, we decided to examine putting an RF Amp on the PSL table, between the RF combiner and the triple resonant box.

5) No proper stuff from Teledyne Couger

 

By looking at what Daniel used in the low noise EOM Driver for aLIGO, we found the A2CP2596 from Cougar.

G = +24 dB, NF = 5 dB, Max Out = +37 dBm. It comes in a 2-stage SMA connector package. I've asked Steve to order 2 of them with the appropriate heatsinks.

  5371   Fri Sep 9 16:16:59 2011 kiwamuUpdateIOOspot positions on MC mirrors (in-vac)

The spot positions on the MC mirrors were measured in the vacuum condition.

The obtained spot positions are quite bad and roughly at 2-3 mm level. We have to realign the beam axis and the MC mirrors.

 

    Feb 26 2011      May 08 2011 Aug 2 2011  Aug 10 2011 (in air) Aug 14 2011 (in air)

Aug 23 2011 (in air)

[NEW]  Sep 9 2011
MC1 pit [mm]   1.6   1.9  1.93 -0.858 -0.2 0.1354  -1.55
MC2 pit [mm]   6.4   9.0 9.03 -0.844 -0.8 -0.2522  -2.28
MC3 pit [mm]   1.4   2.0 2.01 -1.03 -0.1 -0.1383  -1.8
MC1 yaw [mm]   -1.5   -1.7 -1.72 -0.847 -1.1 -1.0893  -0.02
MC2 yaw [mm]   1.0   0.2 0.178 0.582 0.6 0.7122  3.52
MC3 yaw [mm]   -1.3   -1.9 -1.87 -1.06 -1.1 -1.5587  -2.145

 

  5370   Fri Sep 9 14:55:03 2011 kiwamuUpdateSUSOSEM pictures on ITMs

The OSEM pictures taken in Sep/6 have been uploaded to Picasa.

https://picasaweb.google.com/foteee

  5369   Fri Sep 9 13:29:09 2011 kiwamuUpdateIOOa history of PZT mirror

Something happened about 8 years ago.

Old iLog entry by AJW (2003/Sep/8)

Old iLog entry by AJW (2003/Sep/9)

Quote:

Last night I noticed that PZT1 didn't work properly

 

  5368   Fri Sep 9 11:59:58 2011 kiwamuUpdateIOOPZT1 doesn't work

Last night I noticed that PZT1 didn't work properly

I am not sure what is going on. Today I will try localizing the cause of the problem.

As far as I remember it was perfectly working at the time just after we readjusted the OSEMs on MC1 and MC3 (Aug 23th)

 

The symptoms are :

  + No response to both pitch and yaw control from EPICS (i.e. C1:LSC-PZT1_X and C1:LSC-PZT1_Y)

  + When a big value (-3 or so) from EPICS was applied, the PZT1 mirror suddenly jumped.

     However it turned out it just corresponded to a state where OOR (Out Of Range) LED lights up.

 

I did some brief checks :

  + checked the voltage going into the HV amplifiers' "MOD" input. Those are the voltage coming out from DACs and controlled from EPICS.

   --> looked healthy. They went from -10 to 10 V as expected (although the HV amp takes up to only +/-5V).

  + swapped the ''MOD" input cables such that C1:LSC-PZT1 controls the PZT2 HV and vice versa.

    --> The PZT2 mirror was still controlable, but the PZT1 mirror still didn't move. So the DAC and EPICS are innocent.

  + swapped the D-dub cables, which are directly going into the feedthroughs, such that the PZT1 HV drives the PZT2 mirrors and vice versa.

    --> the PZT2 mirror became unable to be controlled. For the PZT1 mirror, only PITCH worked smoothly.


  5367   Thu Sep 8 20:13:24 2011 kiwamuUpdateIOOMC is back to full power

[Suresh / Kiwamu]

 The attenuator was removed and now the MC is happily locked with the full power of 1.2 W.

 

(what we did)

 + replaced the perfect reflector, which was before the MCREFL_PD, by a 10% beam splitter like it used to be.

 + removed the attenuator (combination of HWP and PBS).

 + realigned the beam path on the AP table, including the MCREFL path and WFS path.

 + made the aperture of the MC2F camera narrower in order to avoid a saturation.

 + aligned the MC suspensions so that it resonates with the TEM00 mode.

 + put a ND filter on the AS camera

 

(notes)

C1:IOO-MC_RFPD_DCMON = 0.98 (locked)

C1:IOO-MC_TRANS_SUM = 17500 (locekd)

 

(next things to do)

 + measurement of the spot positions on each MC mirror.

 + centering of the beam spot by steering the input mirrors on the PSL table

  5366   Thu Sep 8 13:20:07 2011 steveUpdatePEMvac envelope at atm for 36 days

We tried not to open chambers above 10,000 particles of 0.5 micron cf/min

New items going in:                       2 rasor beam traps, 5 badly oxidized old silver plated  setscrew with spring loaded tips......to be replaced in the future, viton tips for eq screws....some are lose, gold plated allen wrench installed at ITMX bottom, reglued magnet on ITMX

Bad hardware things found:          nylon ball "locking elements" on OSEM locking set screws with screwdriver  slot, lose  1064 nm filter on OSEM pd

Attachment 1: vent70.jpg
vent70.jpg
  5365   Thu Sep 8 10:50:44 2011 steveUpdateVACpumpdown #71 comleted

Vacuum status:  " vacuum normal"   CC1 pressure 5e-5 Torr

Slow roughing through butterfly valve took 11 hours at 1-2 Torr/min speed.  We had a leak at the metal hose connecting o-ring  that became visible  at 200 Torr.

The vacuum envelope pressure rose to 500 mTorr from 450 mTorr  This is 50 mTorr leak plus outgassing over 11 hrs  over night stop =  5  mTorr /  hr

Atm2, Maglev pumpdown details. P2 and P3  are in the foreline of the Maglev. The pressure peaks at 4 Torr beacause the 3 drypumps have  little pumpimg speed at this range.

            External fan is turned on during this one hour period to prevent  Maglev turbo from overheating.

            Same things are  happenning with the two small TP2 and TP3 pumps under load. Their foreline pressure peak at 1 Torr.  They slow down to 45 K RPM from 50K RPM  and warm up to 36-39 C

Precondition: 36 days at atm

Attachment 1: pd#71.jpg
pd#71.jpg
Attachment 2: maglevpd.jpg
maglevpd.jpg
  5364   Wed Sep 7 22:17:04 2011 ranaUpdateIOORF Amp for EOM on PSL Table

After Steve pointed out the 'deep hoop' issue, we decided to examine putting an RF Amp on the PSL table, between the RF combiner and the triple resonant box.

This will reduce the chances of standing waves in the cables and reduce the radiation induced pick-up in the RF PD and Demod electronics.

We would like to send ~10 dBm from the distribution box to the combiner. We also want to able to get as much as ~33 dBm of drive at 11 and 55 MHz. So the amp should have a gain of ~20-30 dB and an operating range of 10-100 MHz.

Also desirable are low distortion (high IP3) and good reverse isolation ( > 40 dB).

Some possibilities so far (please add your RF Google Results here):

1) Mini-Circuits ZHL-1-2W-S:  G = +32 dB, Max Out = +33 dBm, NF = 6 dB, Directivity = 25 dB

2) Mini-Circuits TIA-1000-1R8:  G=+40 dB, Max Out = +36 dBm, NF = 15 dB   (AC Powered, Inst. Amp), Directivity = 58 dB.

3) Mini-Circuits ZHL-2-8: G = +27dB, Max out = +29 dBm, NF = 6dB, Directivity = 32 dB

4) RFbay MPA-10-40: G = +40dB, Max Out = + 30 dBm, NF = 3.3 dB, Rev Iso = 23 dB

5) No proper stuff from Teledyne Couger

 

  5363   Wed Sep 7 21:45:16 2011 kiwamuUpdateVACpump stopped. Pressure is at 450 mtorr

I stopped puming at 9:30 PM and the pressure in P1 is at 450 mtorr.

I followed exactly he procedure that Steve noted on his elog entry.

Here is a plot of the histroy of our pumping today.

vacuum.png

Quote from #5361

 1, close V3

2, close RV1 with torque wheel

3, turn off PR1 & 3

4, disconnect metal hose between RV1 and PR3

 I will start the Maglev tomorrow morning.

  5362   Wed Sep 7 20:44:11 2011 SureshUpdateelogrestarted

Elog crashed / dormant for long time.  A look at the log file indicated that it was busy generating png thumbnails for pdf files.

Restarted at Wed Sep  7 20:41:13 PDT 2011

 

  5361   Wed Sep 7 18:19:37 2011 steveHowToVACstop pump down for overnight

Quote:

Jamie and Steve

We closed ITMX and ITMY chambers and started pumping around 11am

What we did before:

1, turned off AC power  to PZT Jena HV ps

2, checked jam nut positions

3, cheched  single o-ring shims

4, closed psl out shutter

 We are at 30 Torr of 7 hours of pumping with 2 roughing pumps.

Kiwamu will take over the rest of the roughing today. He will keep an eye on  the pumping speed to be  ~1-2 Torr/min  and open up the manual RV1 valve if needed.

The present status is #3023 of  "chamber open to vacuum open" mode and waiting the P1 pressure to drop to 500 mTorr

He will do the following to stop pumping at P1 = 500 mTorr

1, close V3

2, close RV1 with torque wheel

3, turn off PR1 & 3

4, disconnect metal hose between RV1 and PR3

 I will start the Maglev tomorrow morning.

  5360   Wed Sep 7 17:28:41 2011 kiwamuUpdateLSCX arm loss measurement : not good

The measurement itself wasn't good.

I looked at the full 2 kHz data which was taken during the time when I was running the arm loss script on the X arm.

The plot below shows the raw data. The X arm was locked and unlocked sequentially several times.

The ASDC power didn't show a significant difference between the state where it is locked and unlocked.

I am not sure why, but It could be because of a misalginment or some kind of mode-mismatching, which can decrease the coupling efficiency of light going into the cavity.

 

Xarmlossmeasurement.png

 

(some notes)

The raw data were analyzed.

I split the ASDC data into two data, (1) low power state, when the cavity is locked (2) high power state, when the cavity is unlocked.

Then each state was averaged to estimate the averaged ASDC power in each case.

The number I obtained are :

    ASDC when X arm was locked = 54.77755 cnts

    ASDC when X arm was unlocked = 55.45830 cnts

Those numbers correspond to a round trip loss of  78.780778 ppm, which sounds too small for me.

Quote from #5359

To see what was going on I will look at the trend data.
 

 

  5359   Wed Sep 7 16:21:35 2011 kiwamuUpdateLSCarm loss measurement : resluts

Here are the results of the arm loss measurements, which I have done before the vent.

I ran the existing matlab script, called 'armLoss.m', to estimate the loss. The script resides in /scripts/LSC.


(Y arm)

 Round trip loss =  154.668624 +/- 11.343204 ppm

Yarmloss.png

The figure above is a time series of the measurement.

In the lower plot the power in the ASDC_PD are plotted. The green dotted-curve is the power when the Y arm is unlocked.

The blue dotted-curve is the one when the Y arm is locked.

In the upper plot the estimated loss from each combination of locked/unlocked power are plotted.

 


(X arm)

Round trip loss = ????? 50 ppm ?????

The obtained time series looked wired because difference in the ASDC power when the arm was locked/unlocked were small.

This small difference results in such a small loss.

To see what was going on I will look at the trend data.

Xarmloss.png

Quote from #5077

I did the measurement of the arm loss on both X and Y arm by running the armLoss script.

The results will be posted later.

 

  5358   Wed Sep 7 13:28:25 2011 steveSummaryCamerasAll Camera setups a need upgrading

Quote:

I just tried to adjust the ETMY camera and its not very user friendly = NEEDS FIXING.

* Camera view is upside down.

* Camera lens is contacting the lexan viewport cover; this means the focus cannot be adjusted without misaligning the camera.

* There's no strain relief of the camera cables at the can. Needs a rubber cable grommet too.

* There's a BNC "T" in the cable line.

Probably similar issues with some of the other setups; they've had aluminum foil covers for too long. We'll have a camera committee meeting tomorrow to see how to proceed.

 ITMY has been upgraded  here I have the new lenses on hand to do the others when it fit into the schedule.

  5357   Wed Sep 7 12:03:49 2011 steveUpdateVACPump down has started

Jamie and Steve

We closed ITMX and ITMY chambers and started pumping around 11am

What we did before:

1, turned off AC power  to PZT Jena HV ps

2, checked jam nut positions

3, cheched  single o-ring shims

4, closed psl out shutter

  5356   Wed Sep 7 09:21:57 2011 jamieUpdateSUSSUS spectra before close up

Here are all suspension diagonalization spectra before close up. Notes:

  • TMX looks the worst, but I think we can live with it. The large glitch in the UL sensor at around 999423150 (#5355) is worrying. However, it seemed to recover. The spectra below were taken from data before the glitch.
  • ITMY has a lot of imaginary components. We previously found that this was due to a problem with one of it's whitening filters (#5288). I assume we're seeing the same issue here.
  • SRM needs a little more data to be able to distinguish the POS and SIDE peaks, but otherwise it looks ok.
ITMX ITMX.png        pit     yaw     pos     side    butt
UL    0.355   0.539   0.976  -0.500   0.182 
UR    0.833  -1.406  -0.307  -0.118   0.537 
LR   -1.167   0.055   0.717  -0.445   0.286 
LL   -1.645   2.000   2.000  -0.828  -2.995 
SD   -0.747   0.828   2.483   1.000  -1.637 
8.01148
ITMY  ITMY.png        pit     yaw     pos     side    butt
UL    1.003   0.577   1.142  -0.038   0.954  
UR    0.582  -1.423   0.931  -0.013  -1.031  
LR   -1.418  -0.545   0.858   0.008   1.081  
LL   -0.997   1.455   1.069  -0.017  -0.934  
SD   -0.638   0.797   1.246   1.000   0.264
 4.46659
BS  BS.png        pit     yaw     pos     side    butt
UL    1.612   0.656   0.406   0.277   1.031  
UR    0.176  -1.344   1.683  -0.058  -0.931  
LR   -1.824  -0.187   1.594  -0.086   0.951  
LL   -0.388   1.813   0.317   0.249  -1.087  
SD    0.740   0.301  -3.354   1.000   0.035 
 5.49597
PRM  PRM.png        pit     yaw     pos     side    butt
UL    0.546   1.436   1.862  -0.345   0.866  
UR    1.350  -0.564   0.551  -0.055  -0.878  
LR   -0.650  -0.977   0.138   0.023   0.858  
LL   -1.454   1.023   1.449  -0.268  -1.398  
SD    0.634  -0.620  -0.729   1.000   0.611
 5.78216
SRM      
ETMX ETMX.png        pit     yaw     pos     side    butt
UL    0.863   1.559   1.572   0.004   1.029  
UR    0.127  -0.441   1.869   0.480  -1.162  
LR   -1.873  -0.440   0.428   0.493   0.939  
LL   -1.137   1.560   0.131   0.017  -0.871  
SD    1.838   3.447  -0.864   1.000  -0.135 
 5.5259
ETMY  ETMY.png        pit     yaw     pos     side    butt
UL   -0.337   1.275   1.464  -0.024   0.929  
UR    1.014  -0.725   1.414  -0.055  -1.102  
LR   -0.649  -1.363   0.536  -0.039   0.750  
LL   -2.000   0.637   0.586  -0.007  -1.220  
SD    0.057  -0.016   1.202   1.000   0.142 
 4.22572
MC1  MC1.png        pit     yaw     pos     side    butt
UL    0.858   0.974   0.128   0.053  -0.000  
UR    0.184  -0.763   0.911   0.018   0.001  
LR   -1.816  -2.000   1.872   0.002   3.999  
LL   -1.142  -0.263   1.089   0.037   0.001  
SD    0.040   0.036  -0.216   1.000  -0.002 
 5.36332
MC2  MC2.png        pit     yaw     pos     side    butt
UL    1.047   0.764   1.028   0.124   0.948  
UR    0.644  -1.236   1.092  -0.088  -0.949  
LR   -1.356  -0.680   0.972  -0.096   1.007  
LL   -0.953   1.320   0.908   0.117  -1.095  
SD   -0.092  -0.145  -0.787   1.000  -0.065 
 4.029
MC3  MC3.png        pit     yaw     pos     side    butt
UL    1.599   0.343   1.148   0.168   1.101  
UR    0.031  -1.647   1.139   0.202  -1.010  
LR   -1.969   0.010   0.852   0.111   0.893  
LL   -0.401   2.000   0.861   0.077  -0.995  
SD   -0.414   0.392  -1.677   1.000   0.018 
3.61734

 

  5355   Wed Sep 7 08:14:01 2011 steveUpdateSUSfinal OSEM check

All fine, except ITMX_sensor_UL's  60  counts deep hoop for an hour.

Attachment 1: finalcheck.jpg
finalcheck.jpg
Attachment 2: ITMX10min.jpg
ITMX10min.jpg
Attachment 3: finalsum.png
finalsum.png
  5354   Wed Sep 7 00:47:51 2011 JenneUpdateVACPUMP is a GO!

Steve and Jamie:  After Jamie checks the ITM free swingings, please put on the ITM heavy doors and start the pump!  For real this time!!! Yeah!

  5353   Wed Sep 7 00:44:51 2011 JenneUpdateSUSFreeswing all

I just started a freeswing all, as a final check before we pump:


Wed Sep  7 00:43:21 PDT 2011
999416616

Wed Sep  7 00:43:32 PDT 2011
WATCHDOGS WILL BE RESET 5 HOURS AFTER THIS TIME
sleeping for 5 hours...

Jamie: Please do a quickie analysis (at least for the ITMs) before helping Steve with the heavy doors.

I closed the PSL shutter.

Both ITM chambers were checked for tools, so there should be nothing left to do but put the heavy doors on, and begin pumping.

  5352   Wed Sep 7 00:39:34 2011 ranaUpdateSUSITMX adjustments

(What we did)

* Moved SUS to edge of table for OSEM adjustment.

* Leveled the table in this temporary tower position.

* Rotated all OSEMs to give some seperation between magnets and LED/PD packages.

* Moved the upper OSEM bracket a little bit upward.

* All the OSEM holding set screws were short with flat heads; this is annoying since we would like to use them more like thumbscrews. Steve took the long set-screws out of the old ITMX cage and we swapped them. Need to order ~100 silver-plated socket head spare/replacements.

* Took pictures of OSEMs.

* Moved tower back to old position.

* Releveled the table (added one rectangular weight in the NW corner of the table).

* Find that ITMX OSEMs were a couple 100 micron out of position; we adjusted them in-situ in the final position of the tower, trying not to rotate them. All mean voltages now are within 100 mV of ideal half-light.

* Back/front EQ positions adjusted by the screw method. bottom/top stops adjusted earlier.

* OSEM cables tied down with copper wire.

* Increased the incident power up to 91 mW going into MC to temporarily make the POX beam more visible.

* The POX beam was checked. It was exiting from the chamber and going through about the center of the viewport.

ELOG V3.1.3-